OpenStudy (anonymous):

Sam scored 43, 48, and 42 points on three tests. How many points should he get on the fourth test to get an average score of 50?

OpenStudy (anonymous):

(43+48+42+x) divided by 4 = 50

OpenStudy (anonymous):

multiply each side of the equation by 4

OpenStudy (anonymous):

43+48+42+x=200

OpenStudy (anonymous):

then add like terms

OpenStudy (anonymous):

133+x=200

OpenStudy (anonymous):

then subtract 133 from both sides

OpenStudy (anonymous):

i still lost

OpenStudy (anonymous):

to get x=67

OpenStudy (ali2x2):

\[50=43+48+42+x/4\]